OpenStudy (anonymous):

check my answer!!!!!plz Evaluate each polynomial for x = 3 and for x = −3. −x2 − 3x3 − x4

OpenStudy (anonymous):

x = 3 (-3)² -3(3)³ - 3^4 9 - 27 - 81 = -99 x = -3 -9 + 27 + 81 = 99

OpenStudy (anonymous):

gir.. use a calculator

OpenStudy (anonymous):

wat did u get first

myininaya (myininaya):

-(3)^2-3(3)^3-(3)^4 -9-3(27)-81 -9-81-81 -9-162 -171

myininaya (myininaya):

that is for x=3

myininaya (myininaya):

here i will write it pretty...

OpenStudy (anonymous):

-(-3) - 3(-3)3 - (-3)4 3 + 18 + 24 = 45 -(3)2 - 3(3)3 - (3)4 -6 - 27 -12 = -48 -48, 45

myininaya (myininaya):

\[-(3)^2-3(3)^3-(3)^4=-(9)-3(27)-81=-9-81-81=-90-81=-171\]

myininaya (myininaya):

thast is for x=3

myininaya (myininaya):

x=3, then we have -171

myininaya (myininaya):

\[x=-3\] \[-(-3)^2-3(-3)^3-(-3)^4=-(9)-3(-27)-(81)=-9+81-81=-9+0=-9\]